# GRU: Mitigating the Trade-off between Unlearning and Retention for Large Language Models

### v1

@@ Abstract
Large language model (LLM) unlearning has demonstrated its essential role in removing privacy and copyright-related responses, crucial for their legal and safe applications. However, the pursuit of complete unlearning often comes with substantial costs due to its compromises in their general functionality, leading to a notorious trade-off between unlearning and retention. In examining the update process for unlearning dynamically, we find gradients hold essential information for revealing this trade-off. In particular, we look at the varying relationship between retention performance and directional disparities between gradients during unlearning. It motivates the sculpting of an update mechanism derived from gradients from two sources, i.e., harmful for retention and useful for unlearning. Accordingly, we propose Gradient Rectified Unlearning (GRU), an enhanced unlearning framework controlling the updating gradients in a geometry-focused and optimization-driven manner such that their side impacts on other, unrelated responses can be minimized. Specifically, GRU derives a closed-form solution to project the unlearning gradient onto the orthogonal space of that gradient harmful for retention, ensuring minimal deviation from its original direction under the condition that overall performance is retained. Comprehensive experiments are conducted to demonstrate that GRU, as a general framework, is straightforward to implement and efficiently enhances a range of baseline methods through its adaptable and compatible characteristics. Additionally, experimental results show its broad effectiveness across a diverse set of benchmarks for LLM unlearning.
@@ 1 Introduction
Large language models (LLMs)[1, 2] have revolutionized the learning paradigms towards general-purpose language generation and understanding, enabling new frontiers in linguistic adaptability, knowledge representation, and advanced reasoning. These models typically employ architectures based on multi-head attention decoders with billions of learnable parameters and are trained autoregressively on web-derived datasets containing trillions of tokens [3]. Such substantial scaling equips LLMs to tackle a wide array of complex linguistic tasks, thereby demonstrating remarkable capabilities that span diverse domains such as programming, creative writing, and scientific research [4, 5, 6, 7].
While scaling offers remarkable benefits, it also introduces substantial drawbacks. A primary concern is the propensity of LLMs to memorize data [8, 9], potentially reproducing sensitive messages encountered during its pre-training. It encompasses copyright and privacy-related issues [10, 11], highlighting concerns about the potential misuse of LLMs for illicit activities as well as challenges in safeguarding individual rights [12]. Therefore, LLMs should be used in accordance with ethical and legal principles. To remove these undesirable behaviors, it is essential to conduct regular audits to identify sensitive content and subsequently adjust the embedded knowledge within LLMs by removing them.
As the key technique to achieve this goal, LLM unlearning [11, 13] explores strategies to remove parameterized knowledge targeted to be unlearned directly. One of the foundational methods is gradient ascent (GA) [13], which directly minimizes the log-likelihood for targeted data, thereby reducing their probabilities of being generated to nearly zero. However, GA has notably negative impacts on model responses for other, non-targeted data, spurring subsequent work that regularizes unlearning procedures to retain overall model behaviors [14, 15, 16]. However, there remains an inherent trade-off between unlearning and retention, in which preserving common performance comes at the cost of reducing the effectiveness of unlearning [11, 15]. Consequently, we are motivated to raise a pivotal research question:
@@
We first conduct observational experiments to better understand the model update dynamics during the unlearning process. Specifically, we delve into the fundamental component—model gradients. To do so, we separately compute the gradients of the current model on retain (non-targeted) and unlearning (targeted) data [^1] We adopt the term “retain data” to align with existing literature (e.g., the TOFU benchmark [14]), and also because pairing “retain data” with “unlearning data” improves readability and closely reflects the core theme of our work. In particular, we refer to data not targeted to be unlearned as “retain data,” and data targeted to be unlearned as “unlearning data,” throughout this paper., and measure their directional alignment using cosine similarity. Additionally, we track the corresponding retention performance to clearly illustrate how gradient alignment affects the model’s behavior throughout unlearning. In Figure LABEL:fig:cos_risk, we present two representative pairs of visualizations, illustrating these gradient dynamics and retention performance for the representative unlearning methods GA and Negative Preference Optimization (NPO) [15]. This empirical observation motivated the design of our framework. In the following sections, we further substantiate this motivation through a formal and theoretical analysis.
Figure 4: Conceptual illustration of our unlearning framework. Conventional unlearning methods, such as GA, often suffer from declining retention performance, leading to diminished model utility. GRU mitigates this issue by rectifying the original gradients at each step, ensuring reliable unlearning without compromising retention performance.
@@
To this end, we introduce the Gradient Rectified Unlearning (GRU), a general framework to mitigate the trade-off between unlearning and retention with both optimisation and geometry implications. The key insight of GRU lies in the gradient rectification during model updates: The gradients for unlearning are re-projected onto the orthogonal directions with respect to those that are detrimental to retention, thereby ensuring the overall intact performance under a first-order assumption (cf., Section 3.1). Accordingly, examples illustrating the altered behavior of gradient dynamics are shown in Figure LABEL:fig:cos_risk(b) and (d). The directions that potentially harm retention can be estimated by the gradients from a set of data non-targeted for unlearning, which are readily accessible for many well-established benchmarks [14] or can be directly extracted from pre-trained models [17]. Please refer to Figure 4 for a conceptual illustration of our framework.
@@
We further provide a detailed analysis to comprehend the mechanisms behind GRU. For the goal of retention, we demonstrate that GRU offers enhanced reliability over previous unlearning methods. Therein, accurately estimating the retention direction is key to a successful outcome. For the goal of unlearning, those original methods that possess closer gradient directions (i.e., smaller cosine similarity) to that for retention lead to better effectiveness, thereby allowing the rectified unlearning gradients to maintain a substantial magnitude after adjustment. Hence, a proper choice of fundamental unlearning methods remains equally essential. Additionally, a further consideration is that retain data can often be biased in real scenarios. Developing methods independent of retain data is appealing while maintaining retention performance. With the theoretical insight, we can easily arrive at a variant of GRU, task vector rectified unlearning (TRU), with the same optimization mechanism that can easily adapt to challenging retain-data-free setting.
@@
We conduct comprehensive experiments in a variety of well-established unlearning benchmarks, including TOFU [14], WMDP [18], and MUSE [19]. The integration of our GRU with established baselines demonstrates its effectiveness, achieving powerful unlearning capabilities alongside enhanced retention reliability. We also compare TRU with different competitive methods, and its superiority reveals a promising direction of efforts. These results highlight the generality and substantial potential of our approach in effectively mitigating the trade-off between unlearning and retention.
@@ 2 Preliminaries
We consider a pre-trained LLM that models an autoregressive distribution over sequences of tokens. Specifically, for an input sequence $\boldsymbol{s}=[s_{1},s_{2},\dots,s_{|\boldsymbol{s}|}]$, the probability of the sequence is modeled as the product of conditional probabilities of each token given all preceding tokens:
$p(\boldsymbol{s};\boldsymbol{\theta})=\prod_{i=1}^{|\boldsymbol{s}|}p(s_{i}\mid\boldsymbol{s}_{1:i-1};\boldsymbol{\theta}),$
where $\boldsymbol{\theta}$ denotes model parameters, and $\boldsymbol{s}_{1:i-1}$ represents the subsequence consisting of tokens $s_{1}$ through $s_{i-1}$. $\boldsymbol{\theta}$ is typically learned by minimizing the negative log-likelihood (NLL) loss over a large corpus of web-sourced data $\mathcal{D}_{\mathrm{t}}=\{\boldsymbol{s}^{1},\boldsymbol{s}^{2},\dots,\boldsymbol{s}^{m}\}$ of size $m$, which is given by $-1/m\sum_{\boldsymbol{s}\in\mathcal{D}_{\mathrm{t}}}\log p(\boldsymbol{s};\boldsymbol{\theta})$. Pre-trained LLMs have shown their remarkable capabilities [20]. However, these models also face safety concerns due to their reliance on web-sourced data, potentially leading to privacy breaches [21], copyright infringement [22], and potential misuse [23].
@@ 2.1 LLM Unlearning
These concerns motivate emerging studies of LLM unlearning recently, which aims to effectively remove undesirable data points or entire hazardous domains from original models. Formally, let $\mathcal{D}_{\mathrm{u}}=\{\boldsymbol{s}^{1}_{\mathrm{u}},\boldsymbol{s}^{2}_{\mathrm{u}},\dots,\boldsymbol{s}^{n}_{\mathrm{u}}\}$ represents the unlearning dataset, typically a subset of the training data $\mathcal{D}_{\mathrm{t}}$ where $n\ll m$. The primary objectives of LLM unlearning are twofold [11]: a) Removal: The unlearned model, characterized by parameters $\boldsymbol{\theta}_{\text{u}}$, should eliminate the knowledge associated with $\mathcal{D}_{\mathrm{u}}$, thereby reducing its capacity to recall or reproduce any information targeted to be forgotten. b) Retention: The model should retain its performance on the remaining data $\mathcal{D}_{\mathrm{t}}\setminus\mathcal{D}_{\mathrm{u}}$, ensuring that the capabilities on tasks and data unrelated to the unlearning dataset can be preserved in reliable manner. The objectives of removal and retention are both essential for LLM unlearning, which can be interpreted as a bi-objective learning problem [11, 24].
@@ 2.2 Unlearning Methods
In the following, we present several representative methods for unlearning, each addressing ways to remove or preserve retention performance, while striving to mitigating the trade-off between the two goals.
@@
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})\coloneqq\frac{1}{n}\sum_{\boldsymbol{s}\in\mathcal{D}_{\mathrm{u}}}\log p(\boldsymbol{s};\boldsymbol{\theta}),$ (1)
which directly reduces the probabilities of generating contents resembling $\mathcal{D}_{\mathrm{u}}$ to approach zero, thereby leading to effective knowledge removal. However, due to its extremely large strengths of gradient updates, the resulting GA-unlearned models will suffer from excessive unlearning [11], where the model responses for non-targeted data will also be damaged, and thus it is clear that GA fails at retention performance. This motivates a series of subsequent works in order to improve the retention performance of the resulting models as a result of the study.
@@
$\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta})=-\frac{1}{m^{\prime}}\sum_{\boldsymbol{s}\in\mathcal{D}_{\mathrm{r}}}\log p(\boldsymbol{s};\boldsymbol{\theta}),$ (2)
@@
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\textrm{GD}}(\mathcal{D}_{\mathrm{u}},\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta})\coloneqq\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})+\lambda\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta}),$ (3)
@@
where $\lambda$ is a trade-off hyper-parameter, typically set to 1. However, many previous works [14] reveal that the unlearning term, i.e., $\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})$, tends to dominate the dynamics of gradient updates. Therefore, GD may still strongly impact retention performance negatively.
@@
Negative Preference Optimization (NPO) [15] directly refines the objective of GA to mitigate excessive unlearning, of which the formulation is motivated by direct preference optimization, a well-known preference alignment method [25]. NPO segregates the dis-preferred part from DPO, heuristically employing it as the unlearning objective, following the formulation of
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\textrm{NPO}}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})\coloneqq\frac{1}{n}\sum_{\boldsymbol{s}\in\mathcal{D}_{\mathrm{u}}}\frac{2}{\beta}\log\Big{[}1+\big{(}\frac{p(\boldsymbol{s};\boldsymbol{\theta})}{p(\boldsymbol{s};\boldsymbol{\theta}_{\textrm{org}})}\big{)}^{\beta}\Big{]},$ (4)
@@
where $\beta$ is the inverse temperature and $\boldsymbol{\theta}_{\textrm{org}}$ denotes model parameters before unlearning. The effects of NPO in mitigating excessive unlearning can be understood through its gradients, which are equivalent to GA with extra reweighting [15]. This weighting mechanism pays more attention to data that have small impacts on retention. However, the strength of unlearning for NPO is weaker than that for GA, which could lead to inadequate unlearning.
@@
Unlearning with Control (UWC) [16] suggests a post-unlearning calibration framework. UWC blends model parameters from before and after unlearning to restore retention performance. With a meticulous-searched controlling parameter $\alpha$, we have the calibrated model of
$\alpha\boldsymbol{\theta}_{\textrm{u}}+(1-\alpha)\boldsymbol{\theta}_{\textrm{org}},$ (5)
@@
whose performance on $\mathcal{D}_{\mathrm{t}}\setminus\mathcal{D}_{\mathrm{f}}$ can approach that of $\boldsymbol{\theta}_{\textrm{org}}$. Despite UWC’s flexibility in accommodating a wide range of unlearning methods, its ability to deal with excessive unlearning still comes at the expense of compromises in the effects of unlearning, as it also supports various unlearning methods, alongside its abilities to handle excessive unlearning.

### v3

# GRU: Mitigating the Trade-off between Unlearning and Retention for LLMs

@@ Abstract
Large language model (LLM) unlearning has demonstrated its essential role in removing privacy and copyright-related responses, crucial for their legal and safe applications. However, the pursuit of complete unlearning often comes with substantial costs due to its compromises in their general functionality, leading to a notorious trade-off between unlearning and retention. It motivates this paper to explore enhanced unlearning schemes that can mitigate this trade-off. Specifically, we propose Gradient Rectified Unlearning (GRU), an improved framework that regulates the directions of gradient updates during the unlearning procedure such that their side impacts on other, unrelated responses can be minimized. GRU is easy and general to implement, demonstrating practical effectiveness across a variety of well-established unlearning benchmarks. Our code is available at https://github.com/tmlr-group/GRU.
Machine Learning, ICML
[…]
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Bai et al., 2023; Liu et al., 2024a) have revolutionized the learning paradigms towards general-purpose language generation and understanding. These models employ architectures based on multi-head attention decoders with billions of learnable parameters and are trained autoregressively on web-derived datasets containing trillions of tokens (Brown et al., 2020; Radford et al., 2021; Achiam et al., 2023). Such substantial scaling equips LLMs to tackle a wide array of complex linguistic tasks, showing remarkable capabilities across a diverse range of language tasks (Azerbayev et al., 2023; Roziere et al., 2023; Wu et al., 2023; Thirunavukarasu et al., 2023).
While scaling offers remarkable benefits, it also introduces substantial drawbacks. A primary concern is the propensity of LLMs to memorize data (Petroni et al., 2019; Belrose et al., 2023), potentially reproducing sensitive messages encountered during its pre-training. It encompasses copyright and privacy-related issues (Yao et al., 2023a; Liu et al., 2024b), highlighting concerns about the potential misuse of LLMs for illicit activities as well as challenges in safeguarding individual rights (Zhang et al., 2023). To remove these undesirable behaviors, it is essential to conduct regular audits to identify sensitive content and subsequently adjust the embedded knowledge within LLMs by removing them. This process is crucial for ensuring that the usage of LLMs complies with ethical and legal standards.
As the key technique to achieve this goal, LLM unlearning (Yao et al., 2023b; Liu et al., 2024b; Wang et al., 2024b) explores strategies to directly remove parameterized knowledge targeted to be unlearned. One of the foundational methods is gradient ascent (GA) (Yao et al., 2023b), which directly minimizes the log-likelihood for targeted data, thereby reducing their probabilities of being generated to nearly zero. However, GA has notably negative impacts on model responses for other, non-targeted data, spurring subsequent works that regularize unlearning procedures to retain overall model behaviors (Maini et al., 2024; Zhang et al., 2024; Wang et al., 2024a). Nevertheless, there remains an inherent trade-off between unlearning and retention, in which preserving the common performance comes at the cost of reducing the effectiveness of unlearning (Zhang et al., 2024; Liu et al., 2024b; Wuerkaixi et al., 2025). It motivates us to raise a pivotal research question:
@@
We first conduct observational experiments to better understand the model update dynamics during the unlearning process. Specifically, we delve into the fundamental component—model gradients. To do so, we separately compute the gradients of the current model on retain (non-targeted) and unlearning (targeted) data, and measure their directional alignment using cosine similarity111 We refer to data that are not targeted for unlearning as “retain data” and data targeted to be unlearned as “unlearning data,” aligning with existing literature (Maini et al., 2024). . Additionally, we track the corresponding retention performance to clearly illustrate how gradient alignment affects the model’s behavior throughout unlearning. In Figure LABEL:fig:cos_risk, we present two representative pairs of visualizations, illustrating these gradient dynamics and retention performance for the representative unlearning methods GA and Negative Preference Optimization (NPO) (Zhang et al., 2024). This empirical observation motivated the design of our framework. In the following sections, we further substantiate this motivation through a formal and theoretical analysis.
@@
To this end, we introduce the Gradient Rectified Unlearning (GRU), a general framework to mitigate the trade-off between unlearning and retention with both optimisation and geometry implications. The key insight of GRU lies in the gradient rectification during model updates: The gradients for unlearning are re-projected onto the orthogonal directions with respect to those that are detrimental to retention, thereby ensuring the overall intact performance under a first-order assumption (cf., Section 3.1). Accordingly, examples illustrating the altered behavior of gradient dynamics are shown in Figure LABEL:fig:cos_risk(b) and (d). The directions that potentially harm retention can be estimated by the gradients from a set of data non-targeted for unlearning, which are readily accessible for many well-established benchmarks (Maini et al., 2024) or can be directly extracted from pre-trained models (Carlini et al., 2021). Please refer to Figure 4 for a conceptual illustration of our framework.
@@
We further provide a detailed analysis to comprehend the mechanisms behind GRU. For the goal of retention, we demonstrate that GRU offers enhanced reliability over previous unlearning methods. Therein, an accurate estimation of the retention direction is crucial for its success. For the goal of unlearning, those original methods that possess gradient directions that are closer (i.e., smaller cosine similarity) to that for retention lead to better effectiveness, thereby allowing the rectified unlearning gradients to maintain a substantial magnitude after adjustment. Hence, a proper choice for the basic unlearning methods is also important.
@@
We conduct comprehensive experiments across a variety of well-established unlearning benchmarks, including TOFU (Maini et al., 2024), WMDP (Li et al., 2024), and MUSE (Shi et al., 2024). The integration of our GRU with established baselines demonstrates its effectiveness, achieving powerful unlearning capabilities alongside enhanced retention reliability. These results underscore the generality and significant potential of our approach in effectively mitigating the trade-off between unlearning and retention.
Figure 4: Illustration of our unlearning method. Conventional unlearning methods, such as GA, often suffer from declining retention performance, leading to diminished model utility. GRU mitigates this issue by rectifying the original gradients at each step, ensuring reliable unlearning without compromising retention.
@@ 2 Preliminaries
We consider a pre-trained LLM that models an autoregressive distribution over sequences of tokens. Specifically, for an input sequence $\bm{s}=[s_{1},s_{2},\dots,s_{|\bm{s}|}]$, the probability of the sequence is modeled as the product of conditional probabilities of each token given all preceding tokens:
$p(\bm{s};\bm{\theta})=\prod_{i=1}^{|\bm{s}|}p(s_{i}\mid\bm{s}_{1:i-1};\bm{\theta}),$
where $\bm{\theta}$ denotes model parameters, and $\bm{s}_{1:i-1}$ represents the subsequence consisting of tokens $s_{1}$ through $s_{i-1}$. $\bm{\theta}$ is typically learned by minimizing the negative log-likelihood (NLL) loss over a large corpus of web-sourced data $\mathcal{D}_{\mathrm{t}}=\{\bm{s}^{1},\bm{s}^{2},\dots,\bm{s}^{m}\}$ of size $m$, which is given by $-1/m\sum_{\bm{s}\in\mathcal{D}_{\mathrm{t}}}\log p(\bm{s};\bm{\theta})$. Pre-trained LLMs have shown their remarkable capabilities (Zhao et al., 2023). However, these models also face safety concerns due to their reliance on web-sourced data, potentially leading to privacy breaches (Das et al., 2024), copyright infringement (Eldan & Russinovich, 2023), and potential misuse (Yao et al., 2024).
@@ 2.1 LLM Unlearning
These concerns motivate the emerging studies of LLM unlearning recently, which aims to effectively remove undesirable data points or entire hazardous domains from the original models. Formally speaking, let $\mathcal{D}_{\mathrm{u}}=\{\bm{s}^{1}_{\mathrm{u}},\bm{s}^{2}_{\mathrm{u}},\dots,\bm{s}^{n}_{\mathrm{u}}\}$ represents the unlearning dataset, typically a subset of the training data $\mathcal{D}_{\mathrm{t}}$ where $n\ll m$. The primary objectives of LLM unlearning are twofold (Liu et al., 2024b):
  • a)
    Removal: The unlearned model, characterized by parameters $\bm{\theta}_{\text{u}}$, should eliminate the knowledge associated with $\mathcal{D}_{\mathrm{u}}$, thereby reducing its capacity to recall or reproduce any information targeted to be forgotten.
  • b)
    Retention: The model should also retain its performance on the remaining data $\mathcal{D}_{\mathrm{t}}\setminus\mathcal{D}_{\mathrm{u}}$, ensuring that the capabilities on tasks and data unrelated to the unlearning dataset can be preserved in reliable manner.
The objectives of removal and retention are both essential for LLM unlearning, which can be interpreted as a bi-objective learning problem (Liu et al., 2024b; Wang et al., 2024b).
@@ 2.2 Unlearning Methods
In the following, we present several representative methods for unlearning, each addressing ways to remove or preserve retention performance, while striving to mitigate the trade-off between the two goals. We further discuss gradient projection, a foundational strategy in machine learning, and its recent advances in addressing competing objectives.
@@
$\min_{\bm{\theta}}\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\bm{\theta})\coloneqq\frac{1}{n}\sum_{\bm{s}\in\mathcal{D}_{\mathrm{u}}}\log p(\bm{s};\bm{\theta}),$ (1)
which directly reduces the probabilities of generating contents resembling $\mathcal{D}_{\mathrm{u}}$ to approach zero, thereby leading to effective knowledge removal. However, due to its extremely large strengths of gradient updates, the resulting GA-unlearned models will suffer from excessive unlearning (Liu et al., 2024b), where the model responses for non-targeted data will also be damaged, i.e., GA is not good at retention. It motivates a series of subsequent works to improve the retention performance for the resulting models.
@@
$\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta})=-\frac{1}{m^{\prime}}\sum_{\bm{s}\in\mathcal{D}_{\mathrm{r}}}\log p(\bm{s};\bm{\theta}),$ (2)
@@
$\min_{\bm{\theta}}\mathcal{L}_{\textrm{GD}}$ $(\mathcal{D}_{\mathrm{u}},\mathcal{D}_{\mathrm{r}};\bm{\theta})$ (3)
$\coloneqq\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\bm{\theta})+\lambda\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta}),$
@@
where $\lambda$ is a trade-off hyper-parameter, typically set to 1. However, many previous works (Maini et al., 2024) reveal that the unlearning term, i.e., $\mathcal{L}_{\textrm{GA}}(\mathcal{D}_{\mathrm{u}};\bm{\theta})$, tends to dominate the dynamics of gradient updates. Therefore, GD may still strongly impact retention performance negatively.
@@
Negative Preference Optimization (NPO) (Zhang et al., 2024) directly refines the objective of GA to mitigate excessive unlearning, of which the formulation is motivated by direct preference optimization, a well-known preference alignment method (Rafailov et al., 2024). NPO segregates the dis-preferred part from DPO, heuristically employing it as the unlearning objective, following the formulation of
$\min_{\bm{\theta}}\mathcal{L}_{\textrm{NPO}}$ $(\mathcal{D}_{\mathrm{u}};\bm{\theta})$ (4)
$\coloneqq\frac{1}{n}\sum_{\bm{s}\in\mathcal{D}_{\mathrm{u}}}\frac{2}{\beta}\log\Big{[}1+\big{(}\frac{p(\bm{s};\bm{\theta})}{p(\bm{s};\bm{\theta}_{\textrm{org}})}\big{)}^{\beta}\Big{]},$
@@
where $\beta$ is the inverse temperature and $\bm{\theta}_{\textrm{org}}$ denotes model parameters before unlearning. The effects of NPO in mitigating excessive unlearning can be understood through its gradients, which are equivalent to GA with extra reweighting (Zhang et al., 2024). This weighting mechanism pays more attention to data that have small impacts on retention. However, the strength of unlearning for NPO is weaker than that for GA, which could lead to inadequate unlearning.
@@
Unlearning with Control (UWC) (Wang et al., 2024a) suggests a post-unlearning calibration framework. UWC blends model parameters from before and after unlearning to restore retention performance. With a meticulous-searched controlling parameter $\alpha$, we have the calibrated model of
$\alpha\bm{\theta}_{\textrm{u}}+(1-\alpha)\bm{\theta}_{\textrm{org}},$ (5)
@@
whose performance on $\mathcal{D}_{\mathrm{t}}\setminus\mathcal{D}_{\mathrm{f}}$ can approach that of $\bm{\theta}_{\textrm{org}}$. UWC is flexible in integration with various unlearning methods, while its ability to address excessive unlearning still comes at the cost of compromising the effects of unlearning.
Gradient Rectification for Conflicting Goals.
The idea of modifying gradient directions to resolve conflicts between competing objectives has been explored in various domains, including continual learning and multi-task learning (Lopez-Paz & Ranzato, 2017; Yu et al., 2020). This idea was formalized in continual learning by Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017), which constrains the current task’s update so that it does not increase the loss on past tasks, using a quadratic program to project the update direction into the feasible region defined by gradients of previous tasks. Subsequently, similar geometric principles were adopted in multi-task learning, Gradient Surgery for Multi-Task Learning (PCGrad) (Yu et al., 2020) detects gradient conflicts between tasks and projects each task’s gradient to reduce destructive interference. Our work extends the reach of gradient projection methods, providing a theoretical and practical framework tailored to the specific challenge, i.e., the trade-off between unlearning and retention in LLMs.